# Bidirectional Generative Pre-training for Improving Time Series Representation Learning

### v1

@@ B.1 Dataset Description
Sleep-EDF (Sleep-stage only) dataset for Classification. We leveraged only the sleep stage of the Sleep-EDF dataset for the classification task following existing studies (Eldele et al., 2021; Zhang et al., 2022a). In line with conventional practices, we selected the single EEG channel that captures signals from the Fpz-Cz electrode location. Since the Sleep-EDF dataset labels five sleep stages (i.e., W, REM, N1, N2 and N3), we only focused on EEG signals associated with these sleep stages, resulting in a total of 586.4 million timesteps. This dataset was segmented into 195.5K sequences, each with 3,000 timesteps.
@@
PPG-Dalia Regression. PPG-Dalia dataset (Reiss et al., 2019a) is available from Monash University, UEA&UCR Time Series Regression Archive (Tan et al., 2021). This dataset captures photoplethysmograph (PPG) data for motion compensation and heart rate estimation. Data was collected from 15 subjects engaging in daily life activities using both wrist-worn (Empatica E4) and chest-worn (RespiBAN Professional) devices. All signals were sampled at 700 Hz. The ECG recording serve as a ground truth for heart rate. The PPG and 3D-accelerometer data are used to estimate heart rate, while accounting for motion artefacts. This dataset with a total of 16.6M timesteps was divided into 64,697 sequences, with varied lengths of 256 or 512 timesteps in different dimensions. The pre-training on this PPG-Dalia as well as the PTB-XL dataset was utilized for fine-tuning regression tasks.
@@ B.2 Experiment Details for BiTimelyGPT and Baselines
Pre-training Datasets Sleep-EDF PTB-XL PTB-XL & PPGDalia
Downstream Task Classification Classification Regression
Data Size (Timesteps) 1.2B 109.2M 109.2M & 16.6M
Model Parameters 18M 7.5M 7.5M
@@
Decoder Layers 12 8 8
Heads 8 8 8
Hidden Dim (${\bm{Q}}$, ${\bm{K}}$, ${\bm{V}}$, FF) 320,320,640,640 240,240,480,480 240,240,480,480
@@
Enc-Dec Layers 6 & 6 4 & 4 4 & 4
Encoder Layers 12 8 8
Decoder Layers 12 8 8
Heads 8 8 8
Hidden Dim (${\bm{Q}}$, ${\bm{K}}$, ${\bm{V}}$, FF) 384,384,384,1536 288,288,288,1152 288,288,288,1152

### v3

@@ B.1 Dataset Description
Sleep-EDF (Sleep-stage only) dataset for Classification. We leveraged only the sleep stage of the Sleep-EDF dataset for the classification task following existing studies (Eldele et al., 2021; Zhang et al., 2022b). In line with conventional practices, we selected the single EEG channel that captures signals from the Fpz-Cz electrode location. Since the Sleep-EDF dataset labels five sleep stages (i.e., W, REM, N1, N2 and N3), we only focused on EEG signals associated with these sleep stages, resulting in a total of 586.4 million timesteps. This dataset was segmented into 195.5K sequences, each with 3,000 timesteps.
@@
PPG-Dalia Regression. PPG-Dalia dataset (Reiss et al., 2019) is available from Monash University, UEA&UCR Time Series Regression Archive (Tan et al., 2021). This dataset captures photoplethysmograph (PPG) data for motion compensation and heart rate estimation. Data was collected from 15 subjects engaging in daily life activities using both wrist-worn (Empatica E4) and chest-worn (RespiBAN Professional) devices. All signals were sampled at 700 Hz. The ECG recording serve as a ground truth for heart rate. The PPG and 3D-accelerometer data are used to estimate heart rate, while accounting for motion artefacts. This dataset with a total of 16.6M timesteps was divided into 64,697 sequences, with varied lengths of 256 or 512 timesteps in different dimensions. The pre-training on this PPG-Dalia as well as the PTB-XL dataset was utilized for fine-tuning regression tasks.
Dataset Task Features Timesteps Sequences Length Classes
Sleep-EDF Visualization 7 1.2B 400.7K 3,000 None
Sleep-EDF (sleep) Classification 1 586.4M 195.5K 3,000 5
Epilepsy Classification 1 2.0M 11.5K 178 2
PTB-XL Classification 12 109.2M 21,837 5,000 5
EMG Classification 1 306.0K 204 1,500 3
PPG-Dalia Regression 4 16.6M 64,697 256 & 512 None
IEEEPPG Regression 5 3.1M 3,096 1,000 None
RR Regression 2 31.5M 7,870 4,000 None
HR Regression 2 31.8M 7,949 4,000 None
SpO2 Regression 2 31.8M 7,949 4,000 None
@@
PopHR database. The database hosts a massive amount of longitudinal heterogeneous claim data from the provincial government health insurer in Quebec, Canada (Régie de l’assurance maladie du Québec, RAMQ) on health service use (Shaban-Nejad et al., 2016; Yuan et al., 2018). In total, there are approximately 1.3 million participants in the PopHR database, which represents a randomly sampled 25% of the population in the metropolitan area of Montreal between 1998 and 2014. Cohort membership is maintained dynamically by removing deceased residents and actively enrolling newborns and immigrants.
@@ B.2 Experiment Details for BiTimelyGPT and Baselines
Pre-training Datasets Sleep-EDF PTB-XL PTB-XL&PPGDalia PopHR
Downstream Task Classification Classification Regression Classification
Data Size (Timesteps) 1.2B 109.2M 109.2M & 16.6M 2.2M
Model Parameters 18M 7.5M 7.5M 3M
@@
Decoder Layers 12 8 8 4
Heads 8 8 8 4
Dim (${\bm{Q}}$, ${\bm{K}}$, ${\bm{V}}$, FF) 320,320,640,640 240,240,480,480 240,240,480,480 144,144,288,288
@@
Enc-Dec Layers 6 & 6 4 & 4 4 & 4 4 & 4
Encoder Layers 12 8 8 8
Decoder Layers 12 8 8 8
Heads 8 8 8 4
Dim (${\bm{Q}}$, ${\bm{K}}$, ${\bm{V}}$, FF) 384,384,384,1536 288,288,288,1152 288,288,288,1152 144,144,144,576